# Infrared Thermography in the Tokamak à Configuration Variable

### 12 Mar 2026 v1

@@ 4.2 Boundary conditions and surface layer heat transmission factor
For IR in TCV, a python version of the THEODOR code is used, which solves the differential equation (4.4) implicitly. The three boundary conditions for this THEODOR version are
$\begin{cases}\left.\frac{\partial T}{\partial x}\right|_{\text{side}}=0\\ q_{\perp}(x,d_{tile},t)=0\\ T(x,0,t)=T_{layer}(x,t)-\frac{q_{\perp}(x,0,t-\Delta t)}{\alpha_{top}}\end{cases}$ (4.9)
The first equation means that THEODOR assumes that the temperature derivative at the edges of the heat flux profile is zero. The 2nd equation assumes that the heat flux at the bottom of the tile is zero.
Tile erosion and dust deposition can form surface layers that cause an overestimation of the actual tile temperature [8, 9, 10]. To correct for this effect and avoid false negative heat fluxes, a surface layer heat transmission factor is introduced,
@@
where $k_{layer}$ and $d_{layer}$ are respectively the conductivity and thickness of the layer. The third boundary condition relates the temperature of the surface layer $T_{layer}(x,t)$ (i.e., the temperature actually measured by the IR cameras), to the tile temperature just below the surface layer, $T(x,0,t)$. $q_{\perp}(x,0,t-\Delta t)$ is used since $q_{\perp}(x,0,t)$ is not known at the time where the boundary condition is computed.
@@ 10 Acknowledgements
This work has been carried out within the framework of the EUROfusion Consortium, partially funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). The Swiss contribution to this work has been funded in part by the Swiss State Secretariat for Education, Research and Innovation (SERI). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them.

### 01 Apr 2026 v2

@@ 4.2 Boundary conditions and surface layer heat transmission factor
For IR in TCV, a python version of the THEODOR code is used, which solves the differential equation (4.4) implicitly. Three boundary conditions are used for this THEODOR version. Two of them are
$\begin{cases}\left.\frac{\partial T}{\partial x}\right|_{\text{side}}=0\\ q_{\perp}(x,d_{tile},t)=0\end{cases}$ (4.9)
The first equation means that THEODOR presumes a vanishing temperature derivative at the edges of the heat flux profile. The second equation assumes that the heat flux at the bottom of the tile is zero.
Tile erosion and particle deposition can form surface layers (SLs) that cause an overestimation of the actual tile temperature [8, 9, 10]. In TCV, the main SL components are believed to be made of carbon (wall material), deuterium (injected gas), and boron oxides (due to boronizations). To account for this SL and avoid false negative heat fluxes, a surface layer heat transmission factor is introduced,
@@
where $k_{layer}$ and $d_{layer}$ are respectively the conductivity and thickness of the layer. Differently from previous IDL versions, the Python THEODOR does not use use the boundary condition $T(x,0,t)=T_{layer}(x,t)-\frac{q_{\perp}(x,0,t-\Delta t)}{\alpha_{top}}$ (as it was the case in Eq. (4.11) of [6]). Instead, an actual layer is modeled on top of the tile using $\alpha_{top}$ as input. The SL thickness is assumed to be of $d_{layer}=10\;\mu\text{m}$, although in reality this value could change depending on the tile exposition to the plasma. The SL conductivity is then obtained with $k_{layer}=\alpha_{top}d_{layer}$. The SL mass density times heat capacity, $\rho_{layer}c_{p,layer}$, is supposed to be 100 times smaller than the top of the tile just below the SL, which is then used to estimate the SL diffusivity, $D_{layer}=k_{layer}/(\rho_{layer}c_{p,layer})$.
@@ 10 Acknowledgements
The authors thank Michael Faitsch and Dirk Stieglitz for insightful discussions and continued support with THEODOR. This work has been carried out within the framework of the EUROfusion Consortium, partially funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). The Swiss contribution to this work has been funded in part by the Swiss State Secretariat for Education, Research and Innovation (SERI). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them.